DISCHARGE MODULATION TECHNIQUE FOR DETECTION OF TRANSIENT SPECIES PRODUCED IN PULSED SUPERSONIC JET BY COLOR CENTER LASER SPECTROSCOPY

M. ISHIGURO, T. IMAJO, \underline{K. HARADA}, K. TANAKA and T. TANAKA, Department of Chemistry, Faculty of Science, Kyushu University 33, Hakozaki, Higashi-ku, Fukuoka 812-81, Japan.

A discharge modulation technique has been applied to color center laser spectroscopy for the detection of radical species generated in a pulsed supersonic jet ^{a}. Transitions in the ν_{3} band of the CH_{3} radical rotationally cooled to 23 K were recorded by the phase sensitive detection technique with an improved signal to noise ratio. The sensitivity attained corresponds to 6×10^{−5} absorption.

aM. Ishiguro, T. Imajo, K. Harada, and M. Matsubara, K. Tanaka, and T. Tanaka, Chem. Phys. Lett. 263, 629 (1996).